OpenStudy (anonymous):

I need an explanation about Trigonometric Function, please? Someone from Calculus or higher level should know this. Given that sin u= 5/13 and cos v= -3/5, Both in Quadrant II, find the value of tan(u+v)? I have already known the answer of the tan(u+v), but I need an explanation about these initial answers from the textbook: cos u= -sqrt(1-sin²u)= -12/13 and sin v=sqrt(1-cos²v)= 4/5 How do these -12/13 and 4/5 come from? I checked using calculator according to the given sin u and cos v values, and I got different answers. I will much appreciate it. Thanks!

OpenStudy (amistre64):

pythag identities, and knowing the values of sin and cos for each quadrant

OpenStudy (anonymous):

 cos u = adjacent/ hyp. = -12/13 [ cos is negative in quad 2 ] sin v = opp/ hyp = 4/5